### v4

@@ 5.2 Proving the correctness
We are now in a position to prove the correctness of Algorithm 1. Recall our assumption that the unit disk $\{(x,y):x^{2}+y^{2}\leq 1\}$ is the minimum enclosing disk of $S$. In what follows, let $r>0$ be an input value of the decision algorithm.
First of all, it is clear that whenever Algorithm 1 returns a pair of disks (instead of NO), it is always an $r$-solution for $S$. Therefore, if $r<\mathsf{opt}(S)$ (recall the defenition of $\mathsf{opt}(S)$ in Section 2), Algorithm 1 definitely returns NO. As such, it suffices to consider the case $r\geq\mathsf{opt}(S)$. We show below that the algorithm must return an $r$-solution for $S$.
@@
Proof.
The proof for Condition (1) is somewhat similar to the previous work, e.g., [27]; we provide some details here for completeness. By assumption, $S$ is contained in the unit disk $\{(x,y):x^{2}+y^{2}\leq 1\}$ and is thus $1$-coverable. Depending on whether $r\geq 1$, there are two cases.
  • If $r\geq 1$, we must have $D=D^{\prime}=D\cap D^{\prime}$ as $\{D,D^{\prime}\}$ is a tight $r$-solution. In this case, $S\subseteq D\cap D^{\prime}$ and thus $D\cap D^{\prime}$ intersects $\{(x,y):x^{2}+y^{2}\leq 1\}$. There exists a unit disk $D^{\prime\prime}\subseteq D\cap D^{\prime}$ intersecting $\{(x,y):x^{2}+y^{2}\leq 1\}$, and $D^{\prime\prime}$ in turn contains an axis-parallel unit square $\Box$. Note that $\Box\subseteq D^{\prime\prime}\subseteq\{(x,y):x^{2}+y^{2}\leq 4\}$ and $\Box\cap\mathbb{Z}_{/c}^{2}\neq\emptyset$. Therefore, there exists a point $p\in\{(x,y)\in\mathbb{Z}_{/c}^{2}:x^{2}+y^{2}\leq 4\}\subseteq A$ such that $p\in\Box\subseteq D^{\prime\prime}\subseteq D\cap D^{\prime}$ and thus $\{D,D^{\prime}\}$ is $p$-anchored.
  • If $r<1$, since $\mathsf{dist}(D,D^{\prime})\leq(2-\frac{5}{c})\cdot r$, there exists a disk $D^{\prime\prime}\subseteq D\cap D^{\prime}$ of radius $1/c$, which in turn contains an axis-parallel square $\Box$ of side-length $1/c$. Note that one of $D$ and $D^{\prime}$ must intersect the unit disk $\{(x,y):x^{2}+y^{2}\leq 1\}$, because $S\subseteq D\cup D^{\prime}$. As $r<1$, we know that $\Box\subseteq\{(x,y):x^{2}+y^{2}\leq 4\}$. By the same argument as above, $\Box$ contains some point $p\in\{(x,y)\in\mathbb{Z}_{/c}^{2}:x^{2}+y^{2}\leq 4\}\subseteq A$, and thus $\{D,D^{\prime}\}$ is $p$-anchored.
@@
We claim that neither $D$ nor $D^{\prime}$ contains both $a$ and $b$. Indeed, assume to the contrary this is not the case. Then, $|ab|\leq 2r$. This implies $|uu^{\prime}|\leq 2r$ since $(a,b)$ a farthest pair of $S$ and $u,u^{\prime}\in S$. Consequently, the distance between any two points among $u,u^{\prime},\hat{u}$, and $\hat{u}^{\prime}$ is at most $2r$ since the antipodal points $\hat{u}$ and $\hat{u}^{\prime}$ are contained in $D\cap D^{\prime}$. Note that $\mathsf{ctr}(D)$ is the midpoint of $\overline{u\hat{u}}$. Analogously, $\mathsf{ctr}(D^{\prime})$ is the midpoint of $\overline{u^{\prime}\hat{u}^{\prime}}$. As such, we have $\mathsf{dist}(D,D^{\prime})\leq 3r/2$ (refer to the caption of Figure 6 for a detailed explanation). But this incurs contradiction since $\mathsf{dist}(D,D^{\prime})>(2-\frac{5}{c})\cdot r>3r/2$, with $c=100$. Hence, neither $D$ nor $D^{\prime}$ contains both $a$ and $b$.
@@
Similarly, we can prove $o\in D^{\prime}$. Therefore, $\{D,D^{\prime}\}$ is $o$-anchored. ∎
In what follows, it suffices to focus on the case where neither condition in Observation 14 holds. The failure of condition 1 implies $\mathsf{dist}(D,D^{\prime})>(2-\frac{5}{c})\cdot r$. For the failure of condition 2, we can assume without loss of generality that every point $u\in S\cap\partial D$ has its antipodal point on $\partial D$ outside $D^{\prime}$. We argue below that there exists at least one vector $\vec{\gamma}\in\varGamma$ such that when Algorithm 1 considers $\vec{\gamma}$ in the for-loop, it returns a solution in Line 7.
@@
Choosing $\bm{\vec{\gamma}}$.
The choice of $\vec{\gamma}$ is as follows. Let $\ell$ (resp., $\sigma$) be the line (resp., open segment) connecting the two centers $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$. For any point $p\in\partial D$, we denote by $\hat{p}$ the antipodal point of $p$ on $\partial D$. The line $\ell$ intersects $\partial D$ at two antipodal points $s$ and $\hat{s}$, where $s$ (resp., $\hat{s}$) denotes the one farther (resp., closer) to $\mathsf{ctr}(D^{\prime})$. By Lemma 12, there exist $u,v\in S\cap(\partial D\setminus D^{\prime})$ such that $\overline{uv}\cap(\ell\setminus\sigma)\neq\emptyset$. Note that this implies $\overline{uv}\cap\ell\neq\emptyset$ and $\overline{uv}\cap\sigma=\emptyset$. By our assumption, $u,v,\hat{u},\hat{v}\notin D^{\prime}$. Without loss of generality, we can assume that $|su|\leq|sv|$. See Figure 7. The points $s$ and $\hat{s}$ partition $\partial D$ into two arcs; the $u$-arc refers to the one where $u$ lies on. Since $\overline{uv}\cap\ell\neq\emptyset$, $v$ must lie on the arc other than the $u$-arc, which we call the $v$-arc. (If $u=v=s$, we pick an arbitrary arc as the $u$-arc and let the other one be the $v$-arc.) For each vector $\vec{\gamma}\in\varGamma$, we shoot a ray from $\mathsf{ctr}(D)$ with direction $\vec{\gamma}$. These rays intersect $\partial D$ at $c$ points (which evenly partition $\partial D$ into $c$ arcs). Among these $c$ intersection points, we pick the one on the $u$-arc that is closest to $\hat{s}$. Then we define $\vec{\gamma}\in\varGamma$ as the vector corresponding to this intersection point.
Justifying $\bm{\vec{\gamma}}$.
We show that when considering $\vec{\gamma}$, Algorithm 1 returns a solution. Let $\vec{\eta}\in\mathbb{S}^{1}$ be the unit vector with the same direction as the one originated from $\mathsf{ctr}(D)$ to $\hat{s}$. Note that the angle between $\vec{\eta}$ and $\vec{\gamma}$ is at most $\frac{2\pi}{c}$ by the construction of $\varGamma$ and the choice of $\vec{\gamma}$. By rotating and reflecting the coordinate system, we can assume that $\vec{\gamma}=(1,0)$ and the clockwise angle from $\vec{\gamma}$ to $\vec{\eta}$ at most $\frac{2\pi}{c}$; see Figure 7.
For two points $p,q\in\mathbb{R}^{2}$, we write $p\prec q$ if the $x$-coordinate of $p$ is smaller than that of $q$. The angle of an arc of $\partial D$ is defined as the central angle formed by the arc at $\mathsf{ctr}(D)$. The following observation pertains to relative positions of certain points with respect to $u$ and $v$.
Observation 15.
@@
    $p\notin\mathcal{CR}_{r}(\{u\})$ holds for any point $p\in D^{\prime}\setminus D$ with $p\prec\hat{u}$.
Proof.
We prove these three consitions in order.
@@
The first inequality holds because the distance between $s$ and $u$ is at most $\frac{\pi r}{2}$ (this is in turn because the angle of the subarc within the $u$-arc between $s$ and $u$ is at most $\frac{\pi}{2}$ as already discussed above). The equality in the first line holds because $\vec{\eta}$ is the unit vector directed from $s$ to $\mathsf{ctr}(D)$ and the distance between $s$ and $\mathsf{ctr}(D)$ is $r$. The second inequality holds since $f(\mathsf{ctr}(D))=-\mathsf{dist}(D,D^{\prime})\cdot\langle\vec{\eta},\vec{\gamma}\rangle$ and we have assumed $\mathsf{dist}(D,D^{\prime})>(2-\frac{5}{c})\cdot r$. The last one follows from $\langle\vec{\eta},\vec{\gamma}\rangle\geq 1-\frac{2\pi}{c}$, as the angle between $\vec{\gamma}$ and $\vec{\eta}$ is at most $\frac{2\pi}{c}$. The proves the second condition.
@@
We finally prove Condition (3). Let $p$ be any point in $D^{\prime}\setminus D$ with $p\prec\hat{u}$. Our goal is to prove $p\notin\mathcal{CR}_{r}(\{u\})$. Define $\ell^{\prime}$ to be the perpendicular bisector line of $\sigma$. Let $\ell_{p}$ be the vertical line through $p$. Without loss of generality, we assume that $u$ is above $\ell$. Since $\hat{u}$ is the antipodal point of $u$ on $\partial D$ and $\ell$ contains $\mathsf{ctr}(D)$, $\hat{u}$ is below $\ell$. Also, by our choice of $\vec{\gamma}$, the vector $\vec{\eta}$ is downward, i.e., $\mathsf{ctr}(D)$ is higher than $\mathsf{ctr}(D^{\prime})$ (as illustrated in Figure 7).
Figure 8: Illustration of the proof of Condition (3) of Observation 15.
@@
By the definition of $\ell^{\prime}$, the portion of $D^{\prime}$ to the left of $\ell^{\prime}$ must be inside $D$. Since $p\in D^{\prime}\setminus D$, $p$ must be to the right $\ell^{\prime}$. Recall that $\hat{u}$ is to the left of $\ell^{\prime}$. Since $u$ is above $\ell$ while $\hat{u}$ is below $\ell$, we obtain that the triangle $\Delta p\hat{u}u$ has an obtuse angle at $\hat{u}$; see Figure 8. As a consequence, $|up|>|u\hat{u}|=2r$. Hence, there does not exist a radius-$r$ disk covering both $p$ and $u$, and thus $p\notin\mathcal{CR}_{r}(\{u\})$ holds. ∎
Consider the sets $X,Y,Z$ computed in Lines 3-5 of Algorithm 1 for the vector $\vec{\gamma}$. Our goal is to prove that $Z$ is exactly equal to $S\cap D$ (when viewed as a set), which implies $S\backslash Z\subseteq D^{\prime}$ because $S\subseteq D\cup D^{\prime}$. Note that as long as this is true, both $Z$ and $S\backslash Z$ are $r$-coverable, and thus Algorithm 1 will return a solution in Line 7 when considering $\vec{\gamma}$.
Observation 16.
We have $u\in X\subseteq D$, $\{u,v\}\subseteq Y\subseteq D$, and $Z=S\cap D$.
Proof.
@@
We first prove $u\in X\subseteq D$. Recall that $X$ is the maximal $r$-coverable prefix of $S_{\vec{\gamma}}$. According to Observation 15(2), $p_{i}\notin D^{\prime}$ for all indices $i\leq j$, and therefore $\{p_{1},\dots,p_{j}\}\subseteq D$. Hence, $\{p_{1},\dots,p_{j}\}$ is an $r$-coverable, and thus $\{p_{1},\dots,p_{j}\}\subseteq X$. In particular, $u=p_{j}\in X$. To show $X\subseteq D$, we distinguish two cases: $v\in X$ and $v\notin X$.
@@
  • If $v\in X$, then we have $X\subseteq\mathcal{CR}_{r}(\{u,v\})$ since $u,v\in X$ and $X$ is $r$-coverable. By Lemma 13, $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=D\cap D^{\prime}$. As such, we obtain $X\cap D^{\prime}\subseteq D\cap D^{\prime}$, and consequently, $X\cap(D^{\prime}\setminus D)=\emptyset$. This further implies that $X\subseteq D$ since $X\subseteq D\cup D^{\prime}$.
@@
  • If $v\notin X$, we assume $X\nsubseteq D$ for the sake of contradiction. Then, there exists $p_{i}$ such that $\{p_{1},\dots,p_{i}\}\subseteq X$ is $r$-coverable but $p_{i}$ lies outside $D$ (and thus $p_{i}\in D^{\prime}\setminus D$). Since $v\not\in X$, we have $p_{i}\prec v$. As $v\prec\hat{u}$ by Observation 15(1), $p_{i}\prec\hat{u}$ holds. Because $p_{i}\in D^{\prime}\setminus D$ and $p_{i}\prec\hat{u}$, we further have $p_{i}\notin\mathsf{CR}_{r}(\{u\})$ by Observation 15(3).
@@
    On the other hand, since $X$ is $r$-coverable and $u\in X$, $X\subseteq\mathsf{CR}_{r}(\{u\})$ must hold. As $p_{i}\in X$, we have $p_{i}\in\mathsf{CR}_{r}(\{u\})$. We thus obtain contradiction.
Proving $\bm{\{u,v\}\subseteq Y\subseteq D}$.
@@
We now prove $\{u,v\}\subseteq Y\subseteq D$. Recall that $Y$ is the maximal $r$-coverable prefix of $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$. Since $X$ is $r$-coverable, it follows that $X\subseteq\mathcal{CR}_{r}(X)$, which further implies $X\subseteq Y$. As $u\in X$, we have $u\in Y$.
@@
To show $v\in Y$, we claim that all points of $\{p_{1},\dots,p_{k}=v\}\cap\mathcal{CR}_{r}(X)$ lie in $D$. Assume to the contrary this is not true. Then, there is a point $p_{i}$ in $\{p_{1},\dots,p_{k}=v\}\cap\mathcal{CR}_{r}(X)$ such that $p_{i}\notin D$ (and thus $p_{i}\in D^{\prime}\setminus D$). Thus, either $p_{i}=v$ or $p_{i}\prec v$. Since $v\prec\hat{u}$ by Observation 15(1), we have $p_{i}\prec\hat{u}$. Since $p_{i}\in D^{\prime}\setminus D$ and $p_{i}\prec\hat{u}$, we further have $p_{i}\notin\mathsf{CR}_{r}(\{u\})$ by Observation 15(3). On the other hand, since $u\in X$ and $X$ is $r$-coverable, $\mathcal{CR}_{r}(X)\subseteq\mathcal{CR}_{r}(\{u\})$ holds. As $p_{i}\in\mathcal{CR}_{r}(X)$, we obtain $p_{i}\in\mathcal{CR}_{r}(\{u\})$, a contradiction.
@@
The above claim implies that $\{p_{1},\dots,p_{k}=v\}\cap\mathcal{CR}_{r}(X)$ is $r$-coverable, and thus $v$ must be in $Y$. As $\{u,v\}\subseteq Y$ and $Y$ is $r$-coverable, it follows that $Y\subseteq\mathcal{CR}_{r}(\{u,v\})$. According to Lemma 13, $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=D\cap D^{\prime}$. Therefore, $Y\cap D^{\prime}\subseteq D\cap D^{\prime}$, which leads to $Y\subseteq D$ since $Y\subseteq D\cup D^{\prime}$.
Proving $\bm{Z=S\cap D}$.
@@
We finally prove $Z=S\cap D$, using the fact $\{u,v\}\subseteq Y\subseteq D$. Recall that $Z=S\cap\mathcal{CR}_{r}(Y)$. As $Y\subseteq D$ and $Y$ is $r$-coverable, we have $D\subseteq\mathcal{CR}_{r}(Y)$, which implies $S\cap D\subseteq Z$. It suffices to show $Z\subseteq S\cap D$, or equivalently, $Z\subseteq D$. Observe that $Z=(Z\cap D)\cup(Z\cap D^{\prime})$, since $D\cup D^{\prime}$ contains all points in $Z$. Therefore, we only need to show $Z\cap D^{\prime}\subseteq D$. Since $\{u,v\}\subseteq Y$ and $Y$ is $r$-coverable, we have $\mathcal{CR}_{r}(Y)\subseteq\mathcal{CR}_{r}(\{u,v\})$. By Lemma 13, $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=D\cap D^{\prime}$. It follows that
@@
We then have $Z\cap D^{\prime}\subseteq D$ and thus $Z\subseteq D$. ∎

### v5

@@ 5.2 Proving the correctness
We are now in a position to prove the correctness of Algorithm 1. Recall our assumption that the unit disk $\{(x,y):x^{2}+y^{2}\leq 1\}$ is the minimum enclosing disk of $S$. In what follows, let $r>0$ be the input value of the decision algorithm. Note that it suffices to consider the case $r<1$ since otherwise $S$ is contained in a single disk of radius $r$.
First of all, it is clear that whenever Algorithm 1 returns a pair of disks (instead of NO), it is always an $r$-solution for $S$. Therefore, if $r<\mathsf{opt}(S)$ (recall the definition of $\mathsf{opt}(S)$ in Section 2), Algorithm 1 definitely returns NO. As such, it suffices to consider the case $r\geq\mathsf{opt}(S)$. We show below that the algorithm must return an $r$-solution for $S$.
@@
Proof 5.6.
The proof for Condition (1) is somewhat similar to the previous work, e.g., [Sharir-twocenter]; we provide some details here for completeness.
By assumption, $S$ is contained in the unit disk $\{(x,y):x^{2}+y^{2}\leq 1\}$ and is thus $1$-coverable. Recall that $r<1$. Since $\mathsf{dist}(D,D^{\prime})\leq(2-\frac{5}{c})\cdot r$, $r$ is at least $1/2$. Otherwise, a single disk of radius strictly less than $1$ can cover $S$, which contradicts our assumption that the radius of the minimum enclosing disk of $S$ is $1$. Thus, there exists a disk of radius $1/c$ contained in $D\cap D^{\prime}$, and the disk in turn contains an axis-parallel square $\Box$ of side-length $1/c$. Note that one of $D$ and $D^{\prime}$ must intersect the unit disk $\{(x,y):x^{2}+y^{2}\leq 1\}$ because $S\subseteq D\cup D^{\prime}$. As $r<1$, we know that $\Box\subseteq\{(x,y):x^{2}+y^{2}\leq 4\}$, and thus, $\Box\cap\mathbb{Z}_{/c}^{2}\neq\emptyset$. Therefore, $\Box$ contains some point $p\in\{(x,y)\in\mathbb{Z}_{/c}^{2}:x^{2}+y^{2}\leq 4\}\subseteq A$, and thus $\{D,D^{\prime}\}$ is $p$-anchored.
@@
We claim that neither $D$ nor $D^{\prime}$ contains both $a$ and $b$. Indeed, assume to the contrary this is not the case. Then, $|ab|\leq 2r$. This implies $|uu^{\prime}|\leq 2r$ since $(a,b)$ a farthest pair of $S$ and $u,u^{\prime}\in S$. Consequently, the distance between any two points among $u,u^{\prime},\hat{u}$, and $\hat{u}^{\prime}$ is at most $2r$ since the antipodal points $\hat{u}$ and $\hat{u}^{\prime}$ are contained in $D\cap D^{\prime}$. Note that $\mathsf{ctr}(D)$ is the midpoint of $\overline{u\hat{u}}$. Analogously, $\mathsf{ctr}(D^{\prime})$ is the midpoint of $\overline{u^{\prime}\hat{u}^{\prime}}$. As such, we have $\mathsf{dist}(D,D^{\prime})\leq 3r/2$ (refer to the caption of Figure 6 for a detailed explanation). But this leads to a contradiction since $\mathsf{dist}(D,D^{\prime})>(2-\frac{5}{c})\cdot r>3r/2$, with $c=100$. Hence, neither $D$ nor $D^{\prime}$ contains both $a$ and $b$.
@@
Similarly, we can prove $o\in D^{\prime}$. Therefore, $\{D,D^{\prime}\}$ is $o$-anchored.
In what follows, it suffices to focus on the case where neither condition in Lemma 5.5 holds. The failure of condition 1 implies $\mathsf{dist}(D,D^{\prime})>(2-\frac{5}{c})\cdot r$. For the failure of condition 2, every point $u\in S\cap\partial D$ has its antipodal point on $\partial D$ outside $D^{\prime}$, or this holds for $D^{\prime}$. Without loss of generality, we assume that this is true for $D$. We argue below that there exists at least one vector $\vec{\gamma}\in\varGamma$ such that when Algorithm 1 considers $\vec{\gamma}$ in the for-loop, it returns a solution in Line 1.
@@
Choosing $\boldsymbol{\vec{\gamma}}$.
The choice of $\vec{\gamma}$ is as follows. Let $\ell$ (resp., $\sigma$) be the line (resp., open segment) connecting the two centers $\mathsf{ctr}(D)$ and $\mathsf{ctr}(D^{\prime})$. For any point $p\in\partial D$, we denote by $\hat{p}$ the antipodal point of $p$ on $\partial D$. The line $\ell$ intersects $\partial D$ at two antipodal points $s$ and $\hat{s}$, where $s$ (resp., $\hat{s}$) denotes the one farther (resp., closer) to $\mathsf{ctr}(D^{\prime})$. By Lemma 5.1, there exist $u,v\in S\cap(\partial D\setminus D^{\prime})$ such that $\overline{uv}\cap(\ell\setminus\sigma)\neq\emptyset$. Note that this implies $\overline{uv}\cap\ell\neq\emptyset$ and $\overline{uv}\cap\sigma=\emptyset$. By our assumption, $u,v,\hat{u},\hat{v}\notin D^{\prime}$. Without loss of generality, we can assume that $|su|\leq|sv|$. See Figure 7. The points $s$ and $\hat{s}$ partition $\partial D$ into two arcs; the $u$-arc refers to the one where $u$ lies on. Since $\overline{uv}\cap\ell\neq\emptyset$, $v$ must lie on the arc other than the $u$-arc, which we call the $v$-arc. (If $u=v=s$, we pick an arbitrary arc as the $u$-arc and let the other one be the $v$-arc.) For each vector $\vec{\gamma}\in\varGamma$, we shoot a ray from $\mathsf{ctr}(D)$ with direction $\vec{\gamma}$. These rays intersect $\partial D$ at $c$ points (which evenly partition $\partial D$ into $c$ arcs). Among these $c$ intersection points, we pick the one on the $u$-arc that is closest to $\hat{s}$. Then we define $\vec{\gamma}\in\varGamma$ as the vector corresponding to this intersection point.
Justifying $\boldsymbol{\vec{\gamma}}$.
We show that when considering $\vec{\gamma}$, Algorithm 1 returns a solution. Let $\vec{\eta}\in\mathbb{S}^{1}$ be the unit vector with the same direction as the one originated from $\mathsf{ctr}(D)$ to $\hat{s}$. Note that the angle between $\vec{\eta}$ and $\vec{\gamma}$ is at most $\frac{2\pi}{c}$ by the construction of $\varGamma$ and the choice of $\vec{\gamma}$. By rotating and reflecting the coordinate system, we can assume that $\vec{\gamma}=(1,0)$ and the clockwise angle from $\vec{\gamma}$ to $\vec{\eta}$ is at most $\frac{2\pi}{c}$; see Figure 7.
For two points $p,q\in\mathbb{R}^{2}$, we write $p\prec q$ if the $x$-coordinate of $p$ is smaller than that of $q$. The angle of an arc of $\partial D$ is defined as the central angle formed by the arc at $\mathsf{ctr}(D)$. The following observation pertains to the relative positions of certain points with respect to $u$ and $v$.
Lemma 5.7.
@@
    $|pu|>2r$ for any point $p\in D^{\prime}\setminus D$ with $p\prec\hat{u}$. Note that $|pu|>2r$ implies $p\not\in\mathcal{CR}_{r}(\{u\})$.
Proof 5.8.
We prove these three conditions in order.
@@
The first inequality holds because the distance between $s$ and $u$ is at most $\frac{\pi r}{2}$ (this is in turn because the angle of the subarc within the $u$-arc between $s$ and $u$ is at most $\frac{\pi}{2}$ as already discussed above). The equality in the first line holds because $\vec{\eta}$ is the unit vector directed from $s$ to $\mathsf{ctr}(D)$ and the distance between $s$ and $\mathsf{ctr}(D)$ is $r$. The second inequality holds since $f(\mathsf{ctr}(D))=-\mathsf{dist}(D,D^{\prime})\cdot\langle\vec{\eta},\vec{\gamma}\rangle$ and we have assumed $\mathsf{dist}(D,D^{\prime})>(2-\frac{5}{c})\cdot r$. The last one follows from $\langle\vec{\eta},\vec{\gamma}\rangle\geq 1-\frac{2\pi}{c}$, as the angle between $\vec{\gamma}$ and $\vec{\eta}$ is at most $\frac{2\pi}{c}$. This proves the second condition.
@@
We finally prove Condition (3). Let $p$ be any point in $D^{\prime}\setminus D$ with $p\prec\hat{u}$. Our goal is to prove $|pu|>2r$. Define $\ell^{\prime}$ to be the perpendicular bisector line of $\sigma$. Let $\ell_{p}$ be the vertical line through $p$. By the way, we chose $\vec{\gamma}$, $u$ and the vector $\vec{\gamma}$ originating from $\mathsf{ctr}(D)$ lie on the same side of $\ell$. Recall that we have rotated the coordinate system so that $\vec{\gamma}=(1,0)$. Hence, if $u$ is above $\ell$, then the direction of $\vec{\eta}$ is downward, i.e., $\mathsf{ctr}(D)$ is higher than $\mathsf{ctr}(D^{\prime})$ (see Figure 8(a)). If $u$ is below $\ell$, then $\vec{\eta}$ is upward and $\mathsf{ctr}(D^{\prime})$ is higher than $\mathsf{ctr}(D)$ (see Figure 8(b)). In the following, we only consider the case where $u$ is above $\ell$, since the other case can be proved analogously. Since $\hat{u}$ is the antipodal point of $u$ on $\partial D$ and $\ell$ contains $\mathsf{ctr}(D)$, $\hat{u}$ is below $\ell$.
Figure 8: Illustration of the proof of Condition (3) of Lemma 5.7.
@@
By the definition of $\ell^{\prime}$, the portion of $D^{\prime}$ to the left of $\ell^{\prime}$ must be inside $D$. Since $p\in D^{\prime}\setminus D$, $p$ must be to the right of $\ell^{\prime}$. Recall that $\hat{u}$ is to the left of $\ell^{\prime}$. Since $u$ is above $\ell$ while $\hat{u}$ is below $\ell$, we obtain that the triangle $\Delta p\hat{u}u$ has an obtuse angle at $\hat{u}$; see Figure 8. As a consequence, $|up|>|u\hat{u}|=2r$.
Consider the sets $X,Y,Z$ computed in Lines 1-1 of Algorithm 1 for the vector $\vec{\gamma}$. Our goal is to prove that $Z$ is exactly equal to $S\cap D$ (when viewed as a set), which implies $S\backslash Z\subseteq D^{\prime}$ because $S\subseteq D\cup D^{\prime}$. Note that as long as this is true, both $Z$ and $S\backslash Z$ are $r$-coverable, and thus Algorithm 1 will return a solution in Line 1 when considering $\vec{\gamma}$.
Lemma 5.9.
Let $u$ and $v$ be as above. We have $u\in X\subseteq D$, $\{u,v\}\subseteq Y\subseteq D$, and $Z=S\cap D$.
Proof 5.10.
@@
We first prove $u\in X\subseteq D$. Recall that $X$ is the maximal $r$-coverable prefix of $S_{\vec{\gamma}}$. According to Lemma 5.7(2), $p_{i}\notin D^{\prime}$ for all indices $i\leq j$, and therefore $\{p_{1},\dots,p_{j}\}\subseteq D$. Hence, $\{p_{1},\dots,p_{j}\}$ is an $r$-coverable, and thus $\{p_{1},\dots,p_{j}\}\subseteq X$. In particular, $u=p_{j}\in X$. To show $X\subseteq D$, we distinguish two cases: $v\in X$ and $v\notin X$.
@@
  • If $v\in X$, then we have $X\subseteq\mathcal{CR}_{r}(\{u,v\})$ since $u,v\in X$ and $X$ is $r$-coverable. By Lemma 5.3, $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=D\cap D^{\prime}$. As such, we obtain $X\cap D^{\prime}\subseteq D\cap D^{\prime}$, and consequently, $X\cap(D^{\prime}\setminus D)=\emptyset$. This further implies that $X\subseteq D$ since $X\subseteq D\cup D^{\prime}$.
@@
  • If $v\notin X$, we assume $X\nsubseteq D$ for the sake of contradiction. Then, there exists $p_{i}$ such that $\{p_{1},\dots,p_{i}\}\subseteq X$ is $r$-coverable but $p_{i}$ lies outside $D$ (and thus $p_{i}\in D^{\prime}\setminus D$). Since $v\not\in X$, we have $p_{i}\prec v$. As $v\prec\hat{u}$ by Lemma 5.7(1), $p_{i}\prec\hat{u}$ holds. Because $p_{i}\in D^{\prime}\setminus D$ and $p_{i}\prec\hat{u}$, we further have $p_{i}\notin\mathcal{CR}_{r}(\{u\})$ by Lemma 5.7(3).
@@
    On the other hand, since $X$ is $r$-coverable and $u\in X$, $X\subseteq\mathcal{CR}_{r}(\{u\})$ must hold. As $p_{i}\in X$, we have $p_{i}\in\mathcal{CR}_{r}(\{u\})$. We thus obtain contradiction.
Proving $\boldsymbol{\{u,v\}\subseteq Y\subseteq D}$.
@@
We now prove $\{u,v\}\subseteq Y\subseteq D$. Recall that $Y$ is the maximal $r$-coverable prefix of $S_{\vec{\gamma}}\cap\mathcal{CR}_{r}(X)$. As $X\subseteq D$, $D\subseteq\mathcal{CR}_{r}(X)$ holds. Since $v\in D$, we have $v\in S\cap\mathcal{CR}_{r}(X)$. Furthermore, since $X$ is $r$-coverable, it follows that $X\subseteq\mathcal{CR}_{r}(X)$, which further implies $X\subseteq Y$. As $u\in X$, we have $u\in Y$.
@@
To show $v\in Y$, we claim that all points of $\{p_{1},\dots,p_{k}=v\}\cap\mathcal{CR}_{r}(X)$ lie in $D$. Assume to the contrary this is not true. Then, there is a point $p_{i}$ in $\{p_{1},\dots,p_{k}=v\}\cap\mathcal{CR}_{r}(X)$ such that $p_{i}\notin D$ (and thus $p_{i}\in D^{\prime}\setminus D$). Thus, either $p_{i}=v$ or $p_{i}\prec v$. Since $v\prec\hat{u}$ by Lemma 5.7(1), we have $p_{i}\prec\hat{u}$. Since $p_{i}\in D^{\prime}\setminus D$ and $p_{i}\prec\hat{u}$, we further have $p_{i}\notin\mathcal{CR}_{r}(\{u\})$ by Lemma 5.7(3). On the other hand, since $u\in X$ and $X$ is $r$-coverable, $\mathcal{CR}_{r}(X)\subseteq\mathcal{CR}_{r}(\{u\})$ holds. As $p_{i}\in\mathcal{CR}_{r}(X)$, we obtain $p_{i}\in\mathcal{CR}_{r}(\{u\})$, a contradiction.
@@
The above claim implies that $\{p_{1},\dots,p_{k}=v\}\cap\mathcal{CR}_{r}(X)$ is $r$-coverable, and thus $v$ must be in $Y$. As $\{u,v\}\subseteq Y$ and $Y$ is $r$-coverable, it follows that $Y\subseteq\mathcal{CR}_{r}(\{u,v\})$. According to Lemma 5.3, $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=D\cap D^{\prime}$. Therefore, $Y\cap D^{\prime}\subseteq D\cap D^{\prime}$, which leads to $Y\subseteq D$ since $Y\subseteq D\cup D^{\prime}$.
Proving $\boldsymbol{Z=S\cap D}$.
@@
We finally prove $Z=S\cap D$, using the fact $\{u,v\}\subseteq Y\subseteq D$. Recall that $Z=S\cap\mathcal{CR}_{r}(Y)$. As $Y\subseteq D$ and $Y$ is $r$-coverable, we have $D\subseteq\mathcal{CR}_{r}(Y)$, which implies $S\cap D\subseteq Z$. It suffices to show $Z\subseteq S\cap D$, or equivalently, $Z\subseteq D$. Observe that $Z=(Z\cap D)\cup(Z\cap D^{\prime})$, since $D\cup D^{\prime}$ contains all points in $Z$. Therefore, we only need to show $Z\cap D^{\prime}\subseteq D$. Since $\{u,v\}\subseteq Y$ and $Y$ is $r$-coverable, we have $\mathcal{CR}_{r}(Y)\subseteq\mathcal{CR}_{r}(\{u,v\})$. By Lemma 5.3, $\mathcal{CR}_{r}(\{u,v\})\cap D^{\prime}=D\cap D^{\prime}$. It follows that
@@
We then have $Z\cap D^{\prime}\subseteq D$ and thus $Z\subseteq D$.
This completes the proof.
We have thus completed the full proof of the correctness of Algorithm 1, from which we derive Theorem 1 and Corollary 2.